x^2+14x-5=0

I don't get how to solve any of these type of equation completing the square.

Answer by stanbon(75887)

x^2+14x-5=0
I don't get how to solve any of these type of equations completing the square.
---------------------
x^2+14x + ? = 5 + ?
---
x^2 + 14x + 7^2 = 5 + 7^2
------
(x+7)^2 = 54
x+7 = +-sqrt(54)
x+7 = +-sqrt(9*6)
x+7 = +-3*sqrt(6)
x = -7+3sqrt(6) or x = -7-3sqrt(6)
=================================================
Cheers,
Stan H.